# xMIL: Insightful Explanations for Multiple Instance Learning in Histopathology

### v1

@@ 4.2 Histopathology experiments
Datasets and model training. To evaluate the performance of explanations on real-world histopathology prediction tasks, we considered four diverse datasets of increasing task difficulty covering tumor detection, disease subtyping, and biomarker prediction. These datasets were previously used for benchmarking in multiple studies [12, 33, 42, 62].
@@
We generated patches at 20x magnification and obtained 10,454 $\pm$ 6,236 patches per slide across all datasets (mean $\pm$ std.). Features were extracted using the pre-trained CTransPath [39] foundation model and aggregated using AttnMIL or TransMIL. Additional details regarding the datasets and training procedure are described in Appendix A.3.
We report the mean and standard deviation of the test set AUROC over 5 repetitions in Table 2, indicating high and stable performance levels (>0.93 AUROC) on CAMELYON16 and NSCLC. As expected, we saw lower performance levels for HPV, and a higher variance across repetitions. TP53 mutation could only be predicted with AUROCs below 0.8. In all but one case, TransMIL outperformed AttnMIL, with the largest margin observed in the difficult TP53 dataset. Our results generally align with performances reported in previous works [12, 42, 62]. Notably, our models achieved higher performance levels than reported for additive MIL (0.846 AUROC for CAMELYON16, 0.941 AUROC for NSCLC) [33].
@@
Faithfulness evaluation. As the evidence functions $\mathcal{E}$ of our histopathology datasets are unknown, we resorted to assessing faithfulness, i.e., how accurately explanation scores reflect the model prediction [67, 68]. It can be quantified by progressively excluding instances from the most relevant first (MORF) to the least relevant last and measuring the change in prediction score. The area under the resulting perturbation curve (AUPC) indicates how faithfully the identified ordering of the instances affect the model prediction. The lower the AUPC score, the more faithful the method. We calculated AUPC for correctly classified slides. Further methodological details are provided in Appendix A.4.
In Figure 3, we show the perturbation curves and AUPC boxplots for the patch-dropping experiment for TransMIL and AttnMIL in our four datasets. Additionally, we summarize our results in Table 2. To test the difference in the AUPC values among the baseline explanation methods, we performed paired t-tests between the random baseline vs. all methods and xMIL-LRP vs. all other baselines. The p-values were corrected using the Bonferroni method for multiple comparison correction. All tests resulted in significant differences except for random baseline vs. GxI for CAMELYON16 and attention for HNSC HPV.
Figure 3: Patch dropping results for AttnMIL (bottom) and TransMIL (top). In both plots, the first row depicts the perturbation curves, where the solid lines are the average perturbation curve and the shaded area is the standard error of the mean at each perturbation step. Each boxplot on the second row shows the distribution of AUPC values for all test set slides per explanation methods. In each boxplot, the red line marks the median and the red dot marks the mean. Lower perturbation curves and AUPCs represent higher faithfulness.
@@ 5 Extracting insights from xMIL-LRP heatmaps
Figure 4 presents another correctly classified HPV-positive sample. Here, xMIL-LRP outlines spatially consistent slide regions with clear positive evidence, distinct from regions of negative or mixed evidence (top row). Most notably, the subepithelial mucous glands (bottom row), which are not associated with HPV, are correctly highlighted in blue, unlike in the attention map. Heatmaps of additional examples are provided in Appendix A.5. In Figure 6, xMIL-LRP accurately delineates and distinguishes HPV-positive tumor islands from the surrounding stroma. In this simple case, attention also provides a reasonable explanation. In Figure 7, we display a false positive slide. In this case, xMIL-LRP allowed us to identify that the evidence of HPV-positivity can be attributed to an unusual morphology of an HPV-negative tumor that shares some morphological features usually associated with HPV infection (e.g., smaller tumor cells with hyperchromatic nuclei, dense lymphocyte infiltrates).
@@ Acknowledgements
The results shown here are in part based upon data generated by the TCGA Research Network: https://www.cancer.gov/tcga. This work was in part supported by the German Ministry for Education and Research (BMBF) under Grants 01IS14013A-E, 01GQ1115, 01GQ0850, 01IS18025A, 031L0207D, and 01IS18037A. O.E. received funding from the German Ministry for Education and Research (under refs 01IS18056A and 01IS18025A). K.R.M. was partly supported by the Institute of Information & Communications Technology Planning & Evaluation (IITP) grants funded by the Korea government (MSIT) (No. 2019-0-00079, Artificial Intelligence Graduate School Program, Korea University and No. 2022-0-00984, Development of Artificial Intelligence Technology for Personalized Plug- and-Play Explanation and Verification of Explanation).
@@ A.1 Baseline MIL explanation methods
where $f(\textbf{x}_{k})$ is an instance and $g(X)$ the bag representation. The attention scores $a_{k}$ assign to each patch an attribution with $0\leq a_{k}\leq 1$, and have been used as instance-wise explanation scores [3].
In TransMIL, the attention heads deliver self-attention vectors $\mathbf{A}_{h}^{l}\in\mathbb{R}^{(K+1)\times(K+1)}$ for each head $h$ and Transformer layer $l$, recalling that the first token is the class token. Mean pooling is often used for fusing the self-attention matrices of different heads, i.e., $\mathbf{A}^{l}=\left\langle\mathbf{A}_{h}^{l}\right\rangle_{h}$. The attention scores from the class token to the instance tokens can be used as attribution scores, i.e., $\mathbf{A}^{l}_{(1,1+k)}$. Alternatively, attention rollout has been proposed to summarize the self-attention matrices over layers [60]. For a model with $L$ Transformer layers, attention rollout combines $\{\mathbf{A}^{l}\}_{l=1}^{L}$ as $\tilde{\mathbf{A}}=\prod_{l=1}^{L}\check{\mathbf{A}}^{l}$ where $\check{\mathbf{A}}^{l}=0.5\mathbf{A}^{l}+0.5\mathbf{I}$, with $\mathbf{I}$ being the identity matrix. Then, similar to the layer-wise attention scores, the heatmap is defined as the attention rollout of the class token to the instances, i.e., $\tilde{\mathbf{A}}_{(1,1+k)}$.
@@
A.2 Toy experiments: training and evaluation details
@@
Evidence functions. We define the evidence functions for the three datasets as follows. We write $\textbf{x}_{k}\sim\textbf{n}$ to indicate that instance $k$ represents MNIST number $n$.
@@
  • In the 4-Bags dataset, 8 supports classes 1 and 3 but refutes classes 0 and 2, while 9 supports classes 2 and 3 but refutes classes 0 and 1. Hence, for $\textbf{x}_{k}\sim\textbf{8}$, we define $\epsilon_{k}^{(c)}=1$ for $c\in\{1,3\}$ and $\epsilon_{k}^{(c)}=-1$ for $c\in\{0,2\}$. For $\textbf{x}_{k}\sim\textbf{9}$, we set $\epsilon_{k}^{(c)}=1$ for $c\in\{2,3\}$ and $\epsilon_{k}^{(c)}=-1$ for $c\in\{0,1\}$. In all other cases, $\epsilon_{k}^{(c)}=0$.
@@
  • In Pos-Neg, 4, 6, and 8 instances support class 1 and refute class 0, and vice versa for 5, 7, 9. Hence, we set $\epsilon_{k}^{(1)}=1$ and $\epsilon_{k}^{(0)}=-1$ if $\textbf{x}_{k}\sim\{\textbf{4},\textbf{6},\textbf{8}\}$, $\epsilon_{k}^{(1)}=-1$ and $\epsilon_{k}^{(0)}=1$ if $\textbf{x}_{k}\sim\{\textbf{5},\textbf{7},\textbf{9}\}$, and $\epsilon_{k}^{(c)}=0$ otherwise.
@@
  • In Adjacent Pairs, 4 supports class 1 and refutes class 0 if 3 is also present, but is irrelevant otherwise. That is, for $\textbf{x}_{k}\sim\textbf{4}$, we set $\epsilon_{k}^{(1)}=1$ and $\epsilon_{k}^{(0)}=-1$ if 3 is also in in the bag, and $\epsilon_{k}^{(0)}=0$ otherwise. The evidence scores for the other numbers are defined accordingly.
@@
Evaluation metric. We aim to measure whether an explanation method correctly distinguishes instances with positive, neutral, and negative evidence scores. We separate this into two steps: quantify the separation between positive and non-positive instances, and quantify the separation between negative and non-negative instances. Let $\textbf{e}^{(c)}=[\epsilon^{(c)}_{1},\dots,\epsilon^{(c)}_{K}]$ be the evidence scores for some bag $X=\{\textbf{x}_{1},\dots,\textbf{x}_{K}\}$ and class $c$, and $\textbf{s}^{(c)}=[\hat{\epsilon}^{(c)}_{1},\dots,\hat{\epsilon}^{(c)}_{K}]$ be the class-wise explanation scores from some explanation method. We define $\textbf{e}^{(c)}_{pos}=\min(\textbf{e}^{(c)},\textbf{0})$ and $\textbf{e}^{(c)}_{neg}=\min(-\textbf{e}^{(c)},\textbf{0})$ as the binarized positive and negative evidence, and compute
AUPRC-2 $=\frac{1}{2}\cdot\left(\text{AUPRC}(\textbf{e}^{(c)}_{pos},\textbf{s}^{(c)})+\text{AUPRC}(\textbf{e}^{(c)}_{neg},-\textbf{s}^{(c)})\right).$ (2)
@@
Experimental details. Instead of training end-to-end MIL models, we obtained feature vectors with 512 dimensions for each MNIST image via a ResNet18 model pre-trained on Imagenet from the TorchVision library [71]. For each bag, we first sampled a subset of numbers, where each number was selected with a probability of 0.5, and then randomly drew 30 MNIST feature vectors from this subset. We used 2,000 bags for training, 500 for validation, and 1,000 for testing. We trained AttnMIL and TransMIL models with a learning rate of 0.0001 for a maximum of 1000 and 200 epochs for AttnMIL and TransMIL, respectively. We finally selected the model with the lowest validation loss. We repeated each model training 10 times and report means and standard deviations across repetitions. All experiments were run on single CPUs in less than 5 hours.
A.3 Histopathology experiments: data and training details

### v2

@@ 4.2 Histopathology experiments
Datasets and model training. To evaluate the performance of explanations on real-world histopathology prediction tasks, we considered four diverse datasets of increasing task difficulty covering tumor detection, disease subtyping, and biomarker prediction. These datasets had previously been used for benchmarking in multiple studies [12, 33, 46, 69].
@@
We generated patches at 20x magnification and obtained 10,454 $\pm$ 6,236 patches per slide across all datasets (mean $\pm$ std.). Features were extracted using the pre-trained CTransPath [43] foundation model and aggregated using AttnMIL or TransMIL. [^2] We did not include AddMIL in the real-world experiments, as it is difficult to compare heatmaps from different models without having a ground truth like in the toy experiments (Section 4.1). Also notice that faithfulness evaluations are not applicable, since AddMIL explanations are faithful by design [33]. Additional details regarding the datasets and training procedure are described in Appendix A.4.
We report the mean and standard deviation of the test set AUROC over 5 repetitions in Table 2. In all but one case, TransMIL outperformed AttnMIL, with the largest margin observed in the difficult TP53 dataset. Our results generally align with performances reported in previous works [12, 46, 69].
@@
Faithfulness evaluation. As the evidence functions $\mathcal{E}$ of our histopathology datasets are unknown, we resorted to assessing faithfulness, i.e., how accurately explanation scores reflect the model prediction [74, 75]. The primary goal of the faithfulness experiments is to evaluate the ordering of relevance scores (Property 3 of the evidence function in Definition 3.2). Faithfulness can be quantified by progressively excluding instances from the most relevant first (MORF) to the least relevant last and measuring the change in prediction score. The area under the resulting perturbation curve (AUPC) indicates how faithfully the identified ordering of the instances affects the model prediction. The lower the AUPC score, the more faithful the method. We calculated AUPC for correctly classified slides. Further methodological details are provided in Appendix A.5.
In Figure 3, we show the perturbation curves and AUPC boxplots for the patch-dropping experiment for TransMIL in our four datasets (Figure 4 shows the results for AttnMIL). Additionally, we summarize our results in Table 2. To test the difference in the AUPC values among the baseline explanation methods, we performed paired t-tests between the random baseline vs. all methods and xMIL-LRP vs. all other baselines. The p-values were corrected using the Bonferroni method for multiple comparison correction. All tests resulted in significant differences except for random baseline vs. G×$\times$×I for CAMELYON16 and attention for HNSC HPV.
@@
The results also reflect whether the explanation scores contain meaningful positive/negative evidence for the target class (Property 2 of the evidence function in Definition 3.2): if so, we expect the model’s prediction to flip when all patches supporting the target class are excluded. In Figure 3, the model decision always flips when patches are excluded based on xMIL-LRP scores, whereas other methods show inconsistent results.
@@
Figure 3: Patch dropping results for TransMIL. The first row depicts the perturbation curves, where the solid lines are the average perturbation curve and the shaded area is the standard error of the mean at each perturbation step. Each boxplot on the second row shows the distribution of AUPC values for all test set slides per explanation methods. In each boxplot, the red line marks the median and the red dot marks the mean. Lower perturbation curves and AUPCs represent higher faithfulness.
@@ 5 Extracting insights from xMIL-LRP heatmaps
Heatmaps of additional examples are provided in Appendix A.7. In Figure 6, xMIL-LRP accurately delineates and distinguishes HPV-positive tumor islands from the surrounding stroma. In this simple case, attention also provides a reasonable explanation. Figure 7 presents another correctly classified HPV-positive sample. Here, xMIL-LRP outlines spatially consistent slide regions with clear positive evidence, distinct from regions of negative or mixed evidence (top row). Most notably, the subepithelial mucous glands (bottom row), which are not associated with HPV, are correctly highlighted in blue, unlike in the attention map. In Figure 8, we display a false positive slide. In this case, xMIL-LRP allowed us to identify that the evidence of HPV-positivity can be attributed to an unusual morphology of an HPV-negative tumor that shares some morphological features usually associated with HPV infection (e.g., smaller tumor cells with hyperchromatic nuclei, dense lymphocyte infiltrates).
@@ Acknowledgements
The results shown here are in whole or part based upon data generated by the TCGA Research Network: https://www.cancer.gov/tcga. This work was in part supported by the German Ministry for Education and Research (BMBF) under Grants 01IS14013A-E, 01GQ1115, 01GQ0850, 01IS18025A, 031L0207D, 01IS18037A, and BIFOLD24B and by the Institute of Information & Communications Technology Planning & Evaluation (IITP) grants funded by the Korea government (MSIT) (No. 2019-0-00079, Artificial Intelligence Graduate School Program, Korea University and No. 2022-0-00984, Development of Artificial Intelligence Technology for Personalized Plug-and-Play Explanation and Verification of Explanation).
@@ A.1 Baseline MIL explanation methods
where $f(\mathbf{x}_{k})$ is an instance and $g(X)$ the bag representation. The attention scores $a_{k}$ assign to each patch an attribution with $0\leq a_{k}\leq 1$, and have been used as instance-wise explanation scores [3].
In TransMIL, the attention heads deliver self-attention vectors $\mathbf{A}_{h}^{l}\in\mathbb{R}^{(K+1)\times(K+1)}$ for each head $h$ and Transformer layer $l$, recalling that the first token is the class token. Mean pooling is often used for fusing the self-attention matrices of different heads, i.e., $\mathbf{A}^{l}=\left\langle\mathbf{A}_{h}^{l}\right\rangle_{h}$. The attention scores from the class token to the instance tokens can be used as attribution scores, i.e., $\mathbf{A}^{l}_{(1,2:)}$. Alternatively, attention rollout has been proposed to summarize the self-attention matrices over layers [67]. For a model with $L$ Transformer layers, attention rollout combines $\{\mathbf{A}^{l}\}_{l=1}^{L}$ as $\tilde{\mathbf{A}}=\prod_{l=1}^{L}\check{\mathbf{A}}^{l}$ where $\check{\mathbf{A}}^{l}=0.5\mathbf{A}^{l}+0.5\mathbf{I}$, with $\mathbf{I}$ being the identity matrix. Then, similar to the layer-wise attention scores, the heatmap is defined as the attention rollout of the class token to the instances, i.e., $\tilde{\mathbf{A}}_{(1,2:)}$.
@@
For G×$\times$×I, we can then define the relevance score of the $k$-th instance as $\sum_{d}[\nabla\hat{y}(\mathbf{x}_{k})]_{d}x_{kd}$, with $x_{kd}$ being the $d$-th feature of $\mathbf{x}_{k}$.
Integrated gradients (IG) [66] computes the gradients of the model’s output with respect to the input, integrated over a path from a baseline to the actual input. The baseline is typically set to zero, and so we do. The explanation score of the $k$-th instance is computed as $\sum_{d}\text{IG}(x_{kd})$, where the relevance score of the $d$-th feature of the $k$-th instance $\text{IG}(x_{kd})$ is computed as
[…]
where $f$ is the model and $\mathbf{X}$ is the $K\times D$ feature matrix of the bag (with $K$ being the number of instances and $D$ being the number of features for each instance). We used the implementation of IG available in Captum [77] with the internal batch size set to the number of instances in a bag.
@@
Additive MIL
The idea of additive MIL [33] is to make the MIL model inherently interpretable by designing the bag-level prediction to be a sum of individual instance predictions. Let function $f$ be a feature extractor and $\psi_{m},\psi_{p}$ MLPs. In many cases, particularly for Attention MIL [3], a MIL model $p$ can be written as
$p(X)=\psi_{p}\left(\sum_{k=1}^{K}a_{k}f\left(\mathbf{x}_{k}\right)\right)\quad\text{with}\quad a_{k}=\text{softmax}_{k}(\psi_{m}(X)),$ (4)
where $a_{k}$ is the attention score of instance $k$. For Attention MIL, $\psi_{m}$ is defined as the inner part of the softmax function of Equation 2, and $\psi_{p}$ as prediction head outputting class logits. To obtain an additive model, the authors suggest to instead compute
$p(X)=\sum_{k=1}^{K}\psi_{p}\left(a_{k}f\left(\mathbf{x}_{k}\right)\right).$ (5)
This way, the bag prediction becomes the sum of the individual instance predictions $\psi_{p}\left(a_{k}f\left(\mathbf{x}_{k}\right)\right)$, which can be used as instance explanation scores. These instance logits are proportional to the Shapley values of the instances [33]. In our experiments, we consider the proposed additive variant of Attention MIL (AddMIL).
A.2 Layer-wise Relevance Propagation (LRP)
LRP is a method for explaining neural network predictions by redistributing the output’s relevance back through the network to the input features. The redistribution follows a relevance conservation principle, where the total relevance of each layer is preserved as it propagates backward. If $r_{j}^{(l)}$ denotes the relevance of neuron $j$ in layer $l$, conservation means that $\sum_{j}r_{j}^{(l_{1})}=\sum_{i}r_{i}^{(l_{2})}$ holds for any two layers $l_{1}$ and $l_{2}$. As a general principle, LRP posits
$r_{i}^{(l)}=\sum_{j}\frac{q_{ij}}{\sum_{i^{\prime}}q_{i^{\prime}j}}\cdot r_{j}^{(l+1)},$ (6)
with $q_{ij}$ being the contribution of neuron $i$ of layer $l$ relevance $r_{j}^{(l+1)}$. There are “propagation rules” for various layer types [36, 58] that specify $q_{ij}$ for different setups.
Feed forward neural network. The following generic rule holds for propagating relevance through linear layers followed by ReLU [36]:
$r_{i}^{(l)}=\sum_{j}\frac{a_{j}\rho(w_{ij})}{\epsilon+\sum_{i^{\prime}}a_{i^{\prime}}\rho(w_{i^{\prime}j})}\cdot r_{j}^{(l+1)},$ (7)
where $a_{j}$ is the activation of neuron $j$ in layer $l$, $w_{ij}$ the weight from neuron $i$ of layer $l$ to neuron $j$ of layer $l+1$, $\epsilon$ a stabilizing term to prevent numerical instabilities, and $\rho(w_{ij})$ a modification of the weights of the linear layer. For example, if $\rho(w_{ij})=w_{ij}+\gamma\text{max}(w_{ij},0)$, then Equation 7 is called LRP-$\gamma$ rule. For $\gamma=0$, this equation is called LRP-$\epsilon$ rule.
LayerNorm. Assume $\mathbf{z}_{k}$ is the embedding of the $k$-th token and $\mathbf{y}_{k}=\text{LayerNorm}(\mathbf{z}_{k})$ as:
$\mathbf{y}_{k}=\frac{\mathbf{z}_{k}-\text{E}\{\mathbf{z}\}}{\text{std}\{\mathbf{z}\}+\epsilon},$ (8)
where $\text{E}\{\mathbf{z}\}$ and $\text{std}\{\mathbf{z}\}$ are the expected values and standard deviation of the tokens.
For propagating relevance through LayerNorm, Ali et al. [32] suggested the LN-rule as the following:
$R(z_{kd})=\sum_{j}\frac{z_{kd}(\delta_{kj}-\frac{1}{N})}{\sum_{i}z_{id}(\delta_{ij}-\frac{1}{N})}R(y_{jd}),$ (9)
where $\delta_{kj}=\begin{cases}1,&\text{if }k=j\\ 0,&\text{otherwise}\end{cases}$ and $z_{kd}$ is the $d$-th dimension of $\mathbf{z}_{k}$ and $R(z_{kd})$ is the relevance assigned to it. In practice, LN-rule is implemented by detaching $\text{std}\{\mathbf{z}\}$ and handling it as a constant.
A.3 Toy experiments: Training and evaluation details
@@
Evidence functions. We define the evidence functions for the three datasets as follows. We write $\mathbf{x}_{k}\sim\textbf{n}$ to indicate that instance $k$ represents MNIST number $n$.
@@
  • In the 4-Bags dataset, 8 supports classes 1 and 3 but refutes classes 0 and 2, while 9 supports classes 2 and 3 but refutes classes 0 and 1. Hence, for $\mathbf{x}_{k}\sim\textbf{8}$, we define $\epsilon_{k}^{(c)}=1$ for $c\in\{1,3\}$ and $\epsilon_{k}^{(c)}=-1$ for $c\in\{0,2\}$. For $\mathbf{x}_{k}\sim\textbf{9}$, we set $\epsilon_{k}^{(c)}=1$ for $c\in\{2,3\}$ and $\epsilon_{k}^{(c)}=-1$ for $c\in\{0,1\}$. In all other cases, $\epsilon_{k}^{(c)}=0$.
@@
  • In Pos-Neg, 4, 6, and 8 instances support class 1 and refute class 0, and vice versa for 5, 7, 9. Hence, we set $\epsilon_{k}^{(1)}=1$ and $\epsilon_{k}^{(0)}=-1$ if $\mathbf{x}_{k}\sim\{\textbf{4},\textbf{6},\textbf{8}\}$, $\epsilon_{k}^{(1)}=-1$ and $\epsilon_{k}^{(0)}=1$ if $\mathbf{x}_{k}\sim\{\textbf{5},\textbf{7},\textbf{9}\}$, and $\epsilon_{k}^{(c)}=0$ otherwise.
@@
  • In Adjacent Pairs, 4 supports class 1 and refutes class 0 if 3 is also present, but is irrelevant otherwise. That is, for $\mathbf{x}_{k}\sim\textbf{4}$, we set $\epsilon_{k}^{(1)}=1$ and $\epsilon_{k}^{(0)}=-1$ if 3 is also in in the bag, and $\epsilon_{k}^{(0)}=0$ otherwise. The evidence scores for the other numbers are defined accordingly.
@@
Evaluation metric. We aim to measure whether an explanation method correctly distinguishes instances with positive, neutral, and negative evidence scores. We separate this into two steps: quantify the separation between positive and non-positive instances, and quantify the separation between negative and non-negative instances. Let $\mathbf{e}^{(c)}=[\epsilon^{(c)}_{1},\dots,\epsilon^{(c)}_{K}]$ be the evidence scores for some bag $X=\{\mathbf{x}_{1},\dots,\mathbf{x}_{K}\}$ and class $c$, and $\mathbf{s}^{(c)}=[\hat{\epsilon}^{(c)}_{1},\dots,\hat{\epsilon}^{(c)}_{K}]$ be the class-wise explanation scores from some explanation method. We define $\mathbf{e}^{(c)}_{pos}=\min(\mathbf{e}^{(c)},\mathbf{0})$ and $\mathbf{e}^{(c)}_{neg}=\min(-\mathbf{e}^{(c)},\mathbf{0})$ as the binarized positive and negative evidence, and compute
AUPRC-2 $=\frac{1}{2}\cdot\left(\text{AUPRC}(\mathbf{e}^{(c)}_{pos},\mathbf{s}^{(c)})+\text{AUPRC}(\mathbf{e}^{(c)}_{neg},-\mathbf{s}^{(c)})\right).$ (10)
@@
Experimental details. Instead of training end-to-end MIL models, we obtained feature vectors with 512 dimensions for each MNIST image via a ResNet18 model pre-trained on Imagenet from the TorchVision library [79]. For each bag, we first sampled a subset of numbers, where each number was selected with a probability of 0.5, and then randomly drew 30 MNIST feature vectors from this subset. We used 2,000 bags for training, 500 for validation, and 1,000 for testing. We trained AttnMIL and TransMIL models with a learning rate of 0.0001 for a maximum of 1000 and 200 epochs for AttnMIL and TransMIL, respectively. We finally selected the model with the lowest validation loss. We repeated each model training 30 times and report means and standard deviations across repetitions. Each experiment with its repetitions was run on single CPUs in less than 24 hours, respectively. We used the same setting for training AddMIL, but with an Adam optimizer as in the original paper [33].
A.4 Histopathology experiments: Data and training details